OpenStudy (anonymous):

Which ordered pair is a solution of the equation 5x + y = –23?

OpenStudy (anonymous):

There is no true answer to this, it seems that you can only solve 1 variable, you would need a second equation to define a single solution. but you could go with x= 2, that would give you 10+y = -23 then subtract 10 over, and y would have to be -33. so one answer might be (2,-33)? hope that helps, did you get a second equation with this problem ?

OpenStudy (anonymous):

o sorry its a multiple choice question but i got -4,-3

OpenStudy (anonymous):

im not sure if thats right

OpenStudy (anonymous):

it is, ill explain it

OpenStudy (anonymous):

(-4,-3) means x= -4 and y= -3. so plug those values into the original: 5(-4)+(-3) = -23 then start simplifying it, so -20-3=-23 which you can tell will come out true, therefore (-4,-3) is a solution to the line equation. Does that help?

OpenStudy (anonymous):

yes it helps a whole lot

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

np